OpenStudy (anonymous):

An energy bar company manufactures peanut butter and chocolate bars. Peanut butter bars cost 25¢ per lb. The chocolate bars cost 30¢ per lb. The company sold 260 lbs of bars for $71. How many of the bars sold were peanut butter?

OpenStudy (yanasidlinskiy):

Peanut Butter = P p + c = 260 c = 260 - p .................... Eq. 1 0.25p + 0.3c = 71 25p + 30c = 7100 .......... Eq. 2 Sub 260 - p from Eq. 1 for c in Eq. 2: 25p + 30(260 - p) = 7100 25p + 7800 - 30p = 7100 - 5p = 7100 - 7800 - 5p = - 700 p = - 700 / - 5 p = 140

OpenStudy (nurali):

Let peanut butter bars sold be x lb cost of it 25¢ per lb. = 25x Then chocolate bars sold = (260 – x) cost 30¢ per lb = 30(260 – x) total cost = 25x + 30(260 – x) = 7100 25x + 7800 – 30x = 71 5x = 700 x = 140 -------- peanut butter bars 260 – 140 = 120 chocolate bars